CORRECT ANSWER NEEDED ASAP THROUGH SPREADSHEET

On December 31, 20x0, you make a large sale of inventory to a customer. You

offered your customer financing terms as follows:

• $300,000 payable on December 31, 20x3

• interest of 1.5% on the $300,000 payable on Dec 31, 20x1, 20x2 and

20x3.

Your incremental borrowing rate is 6% and you estimate that your customer’s

incremental borrowing rate is 8%.

Prepare all journal entries for this transaction for the years ended December 31,

20x0, 20x1 and 20x2.
Answer #1
Year PV factor @ 6% Remarks
1    0.94340 = 1 / 1.06
2    0.89000 = 0.9434 / 1.06
3    0.83962 = 0.89 / 1.06
Total    2.67301
Amount Multiply: PV factor Present value
Interest receipts (300000*1.5%)         4,500        2.67301            12,029
Principal    300,000        0.83962          251,886
Value of sales revenue          263,915
effective interest method
Date

Interest recipt

[A]

Interest revenue @6%

[B]

Discount amortized

C = B - A

carrying amount of note

D = Previous' D - C

Year 0          263,915
Year 1         4,500      15,835           11,335          275,250
Year 2         4,500      16,515           12,015          287,265
Year 3         4,500      17,235           12,735          300,000

Interest revenue = Previous period's carrying amount of note * 6%

Date Accounts title Debit Credit
December 31, 20x0 Notes receivable             300,000
Discount on Note receivable (300000-263915)              36,085
Sales revenue            263,915
To record notes received for selling the goods.
December 31, 20x1 Cash                  4,500
Discount on Note receivable                11,335
Interest revenue              15,835
To record interest revenue on notes.
December 31, 20x2 Cash                  4,500
Discount on Note receivable                12,015
Interest revenue              16,515
To record interest revenue on notes.
December 31, 20x3 Cash                  4,500
Discount on Note receivable                12,735
Interest revenue              17,235
To record interest revenue on notes.
December 31, 20x3 Cash             300,000
Notes receivable            300,000
To records cash receipts for principal of notes.

Discount on Note receivable can be replaced by "Unearned interest revenue".